OpenStudy (anonymous):

Given the parent functions f(x)= log2 (3x-9) and g(x)= log2 (x-3). what is f(x)*g(x)

OpenStudy (anonymous):

Sorry I mean f(x)-g(x)

OpenStudy (yanasidlinskiy):

@johnweldon1993 is smart!!! You might want to ask him!:)

OpenStudy (johnweldon1993):

Is that log(base2) ? \[\large log_2\] ?

OpenStudy (anonymous):

Yes it is

OpenStudy (johnweldon1993):

Alright...so remember this rule \[\large log_a (b) - log_a (c) =log_a(\frac{b}{c})\] so really with your function we have \[\large f(x) - g(x)\] \[\large log_2 (3x - 9) - log_2 (x - 3)\] \[\large log_2 (3x - 9) - log_2 (x - 3) \rightarrow log_2 (\frac{3x - 9}{x - 3})\] \[\large log_2 (\frac{3x - 9}{x - 3})\] can you solve that?

OpenStudy (johnweldon1993):

Hint* Factor out a '3' from the numerator

OpenStudy (johnweldon1993):

\[\large \log_2(\frac{3(x - 3)}{(x - 3)})\] \[\large log_2 \frac{3\cancel{(x - 3)}}{\cancel{(x - 3)}}\] \[\large log_2 3 = \space?\]

OpenStudy (anonymous):

@johnweldon1993 you still there?

OpenStudy (anonymous):

Its gonna be 3! :)